OpenStudy (anonymous):

[x^(2n+1)+4x^(2n)-3x^(n+1)]/x^n please somebody help me..

OpenStudy (anonymous):

Can you specify your problem? Such as, do you want us to simply the aforementioned expression?

OpenStudy (anonymous):

the question is like this. perform the indicated operation and simplify.

OpenStudy (anonymous):

I see. [x^(2n+1)+4x^(2n)-3x^(n+1)]/x^n=(x^(2n+1))/(x^n)+(4x^(2n))/(x^n)-(3x^(n+1))/(x^n) =x^(2n+1-n)+4*x^(2n-n)-3*x^(n+1-n) =x^(n+1)+4*x^n-3*x The rest depends on where you want to go from my last step. I hope this helps.

OpenStudy (anonymous):

owhhh... okay2.. i'll try it 1st. sorry for the late-respond

OpenStudy (anonymous):

okay i get it.. thank u sooo much